zombieblud:

I'm not normally the math type of person, but I'm gonna explain Pythagorean theorem to some of those who don't understand it. (pictures will be posted below in comments) a² + b² = c² is what you use to solve for the triangle. To find a² + b² you look at the triangle and take the top and the middle line as a² + b². the bottom of the triangle is c². after you plug in the values for a² + b², you square it. (squaring will be taking a number times itself. ex: 3.14²= 3.14 x 3.14= 9.8596). after you add the squared numbers together you square it again just to get your answer.

zombieblud:

a² + b² = c²

1 attachment
Uniful3:

yay

zombieblud:

I will post how to find C later

BeautyQueen:

So how you find the hypotenuse of a triangle which is C is you use pythagoras theorem Which is a^2 plus b^2 = c^2. So you take the square root of the sum squares and your formula is c = √(a² + b²) which is your formula.